[S] UM 320 160 80.0 40.0 20.0 10.0 Volumol nk-1 S-1) 169 132 92.0 57.2 32.6 17.5

(b.) calculate KM for the enzyme

(c) When the reactions in part (b) are repeated in the presence of 12 microm of an uncompetitive inhibitor, the y-intercept of the Lineweaver-Burk plot is 0.352 (micromol-1 ml s). Calculate K'I for this inhibitor.
Answer #1

Michaelis Menten equation is

v = Vmax / (1 + (Km/[S]))

Lineweaver burk plot is between 1/[S] on x-axis and 1/Vo on y-axis.

Equation of this graph is y = 0.5287x + 0.0043

According to this equation, slope = 0.5287 and intercept = 0.0043

Also, intercept = 1/Vmax = 0.0043

Vmax = 232.55 uM/min

Slope = Km/Vmax = 0.5287

So, Km = 0.5287 *Vmax = 0.5287 *232.55= 122.95 uM

Answer : Km = 122.95 uM

Vmax = 232.55 uM/min

In presence of inhibitor

Vapp = Vmax / (1 + [I]/Ki)

since intercept = 1/vapp = 0.352 hence apparent Vmax = 2.84 uM/min

putting the values 2.84 = 232.55 / (1 + [12 uM] / Ki)

on solving Ki = 0.148361